ROTATIONALLY RESOLVED B̃ ← X̃ ELECTRONIC SPECTRA OF THE ISO-PROPOXY RADICAL AND OTHER SECONDARY ALKOXY RADICALS

\underline{J. LIU}, and T. A. MILLER, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, OH 43210, USA.

The B̃ −X̃ laser-induced-fluorescence (LIF) spectrum of jet-cooled iso-propoxy radical (i-C_{3}H_{7}O·) has been recorded. The observed rotational and fine structure of the origin band has been well simulated with rotational constants for both the X̃ and B̃ states and the electron spin-rotation constants of the X̃ state. The line intensities are well simulated with a parallel transition type, requiring the same symmetry for the X̃ as for the B̃ state (A′), which confirms the previous suggestion^{a} that going from ethoxy (C_{2}H_{5}O·) to iso-propoxy, the energy ordering of the in- and the out-of-plane half-filled p-orbitals is reversed and the ground state symmetry changes from A′′ to A′. The fit rotational constants are consistent with those obtained from the quantum chemistry calculations. However, the fit spin-rotation constants, particularly the nonzero ϵ_{aa} with a-axis perpendicular to the C_{s} plane, can not be explained by second-order perturbation theory. Nevertheless they can be semi-quantitatively explained based on the transferability of the electron spin-molecular rotation tensor^{b} in the axis system defined by the half-filled p-orbital and the pπ-orbital using ethoxy as the reference molecule. The failure of the second-order perturbation theory, as well as that of the pure precession approximation, is due to the strong vibronic interaction between the near-degenerate X̃ and Ã states. The vibrationless levels of these two states are separated by only 68 cm^{−1} as determined in previous disperse-fluorescence study.^{c} The near degeneracy of the X̃ and Ã states of isopropoxy and reversal of the energy ordering going from ethoxy to iso-propoxy have also been demonstrated in quantum chemistry calculations. Based on the result of iso-propoxy, spin-rotation constants of other secondary alkoxy radicals can be predicted and used for the simulation of the B̃ − X̃ rovibronic spectra of these radicals. Specifically, analysis of the high-resolution LIF spectrum of cyclohexoxy (c-C_{6}H_{11}O·)^{d} is ongoing.

aS. C. Foster and T. A. Miller, J. Phys. Chem. 93, 5986 (1989).

bG. Tarczay, S. Gopalakrishnan, and T. A. Miller, J. Mol. Spectrosc. 220, 276 (2003).

cJ. Jin, I. Sioutis, G. Tarczay, S. Gopalakrishnan, A. Bezant, and T. A. Miller, J. Chem. Phys. 121, 11780 (2004).

dL. Zu, J. Liu, G. Tarczay, P. Dupre, and T. A. Miller, J. Chem. Phys. 120, 10579 (2004).